# Classifier-free graph diffusion for molecular property targeting

### v1

@@ 5 Results
The experimental results are presented in Table 1. Novelty and uniqueness scores are not reported as they approached 100% in almost all cases. On the QM9 dataset (top table), we observe that FreeGress outperforms DiGress in almost every variant tested. Interestingly enough, even variants of FreeGress with $p_{uncond}=0$, i.e. those that use the guide without conditional dropout, achieve comparable performance with DiGress, with the only exception of the dipole moment $\mu$. Overall, the best performing variant of FreeGress ($p_{uncond}=0.1$, $s=5$, additional features) improves with respect of DiGress by 50% on the HOMO task, by 33% on the $\mu$ task, and by 39% on the combination of both. Similar improvements are achieved with slightly different combinations of the hyper-parameters, showing that CF guidance is robustly a better alternative to CB guidance for conditioned generation. Also, one should take into account that CF guidance does not require to train a separate auxiliary regressor, which saves up computational and memory resources. Four conditioned samples (two for each property targeting task) generated from the QM9 dataset are shown in Figure 2.
The analysis of performance in the ZINC-250k provides some additional interesting insights. Firstly, we observe a drop in validity of DiGress whenever too much weight is put on the auxiliary regressor by increasing $\lambda$. This is a clear signal that the gradient of the regressor is not providing useful directions in chemical space, which validates our belief that CB guidance, as formulated in DiGress, is based on unrealistic assumptions for the chemical domain. Secondly, we observe that FreeGress outperforms again DiGress on the two tasks by a great margin. Specifically, on the LogP task, FreeGress achieves a 69% improvement in MAE with respect to the best DiGress variant, while at the same time improving validity by 31%. On the MW task, the improvement in MAE is of about 79% with respect to the best DiGress variant. Thirdly, we observe that node inference has not a decisive impact in the LogP task. However, in the MW task, it improves the validity rate from 58% to 75% (on par with the best DiGress variant), while also slightly improving in terms of MAE. This aligns with our intuition: in fact, molecules with higher molecular weight are in general more likely to have more atoms than molecules with lower molecular weight. Thus, it makes sense, in this particular task, to having a mechanism to infer the number of nodes from the guide to bias the sampling process. Figure 2 shows 4 exemplay molecules (2 for each task) generated from the ZINC-250k dataset.
@@ 6 Conclusions
This work has developed and implemented FreeGress, a discrete DDPM for graphs with CF guidance, which allows to generate molecules complying to a pre-specified set of desired chemical properties. Through experiments, we have shown that CF guidance allows to generate better (more tailored to the specification) molecules than CB guidance without sacrificing chemical validity. Of course, there are still limitations to overcome, which we intend to address in future studies: for example, even if CF guidance allows to achieve good validity rates, the model is still far from achieving perfect validity. This might be addressed by designing the forward and reverse processes to use conditional transition matrices. Another way to increase the percentage of valid molecules generated without overfitting may be to train a model to work with molecular fragments instead of atoms, as it would likely reduce the amount of points in the molecular graph where an illegal structure is generated. We also have shown that inferring the number of nodes from the guide might be of help to generate graphs in some settings, especially when there is reason to believe that the guide is correlated with the graph’s size. As future work, we intend to validate this finding also in the more general setting of arbitrary graph generation.
Input LogP: 0.8472Estimated LogP: 0.8590
Input LogP: 2.1313Estimated LogP: 2.0127
Input MW: 401.13Estimated MW: 411.06
Input LogP: 0.8472Estimated LogP: 0.8590
Input LogP: 2.1313Estimated LogP: 2.0127
Input MW: 401.13Estimated MW: 411.06
Input MW: 195.02Estimated MW: 196.18
Figure 3: Curated molecules from the ZINC-250k dataset.

### v2

@@ 5 Results
5.0.1 QM9.
The experimental results on QM9 are presented in Table 1, where we report the MAE and the chemical validity of the generated molecules. Uniqueness scores are not reported as they approached 100% in almost all cases, while the novelty is not a relevant metric on QM9 as the dataset is an enumeration of all the valid molecules satisfying a particular set of constraints. We observe that FreeGress variants, and especially those that leverage additional features, outperform DiGress variants in the vast majority of cases. Interestingly enough, even FreeGress variants with $\rho=0$, i.e. those that use the guide without conditional dropout, are competitive with DiGress. Overall, FreeGress convincingly outperforms DiGress even when the additional features are not used (by 16%, 38% respectively on the $\mu$ and HOMO properties, and by 37% when conditioning on both). When targeting $\mu$+HOMO in conjunction, the gap in performance indicates that FreeGress is better suited for multiple conditioning. As we show in Supplementary Table 1, this might be attributed to the property regressor of DiGress, which struggles to generate molecules conditioned on both properties, while FreeGress does not suffer from this issue. Overall, we observe a slight inverse correlation between MAE and validity, with one tending to decrease as the other improves.
Table 2: Results on the ZINC-250k dataset. First two columns are single conditioning, last column is multiple conditioning on LogP and QED.
LogP QED LogP+QED
MAE $\downarrow$ Val. $\uparrow$ MAE $\downarrow$ Val. $\uparrow$ MAE $\downarrow$ Val. $\uparrow$
Unconditional $1.52\pm 0.12$ $86.1\%$ $0.15\pm 0.01$ $86.1\%$ $0.83\pm 0.06$ $86.1\%$
DiGress
$\lambda=200$ $0.83\pm 0.08$ $77.6\%$ $0.15\pm 0.01$ $85.8\%$ $0.49\pm 0.03$ $83.8\%$
$\lambda=400$ $0.74\pm 0.08$ $74.6\%$ $0.15\pm 0.01$ $85.1\%$ $0.40\pm 0.02$ $78.2\%$
$\lambda=600$ $0.76\pm 0.10$ $69.6\%$ $0.14\pm 0.01$ $84.5\%$ $0.38\pm 0.02$ $79.3\%$
$\lambda=800$ $0.92\pm 0.12$ $65.1\%$ $0.14\pm 0.01$ $83.8\%$ $0.35\pm 0.01$ $75.5\%$
FreeGress
$\rho=0$ $0.22\pm 0.02$ $80.9\%$ $0.06\pm 0.01$ $84.7\%$ $0.15\pm 0.01$ $82.3\%$
$\rho=0.1,s=1$ $0.22\pm 0.02$ $84.7\%$ $0.06\pm 0.01$ $83.8\%$ $0.16\pm 0.01$ $83.1\%$
$\rho=0.1,s=2$ $0.17\pm 0.01$ $84.9\%$ $0.04\pm 0.01$ $84.9\%$ $0.12\pm 0.01$ $80.7\%$
$\rho=0.1,s=3$ $0.18\pm 0.02$ $82.2\%$ $0.05\pm 0.01$ $85.6\%$ $0.12\pm 0.01$ $77.5\%$
$\rho=0.2,s=1$ $0.22\pm 0.02$ $87.0\%$ $0.07\pm 0.01$ $84.2\%$ $0.17\pm 0.02$ $78.6\%$
$\rho=0.2,s=2$ $0.19\pm 0.02$ $83.2\%$ $0.05\pm 0.01$ $82.8\%$ $0.13\pm 0.01$ $76.8\%$
$\rho=0.2,s=3$ $0.16\pm 0.01$ $81.2\%$ $0.05\pm 0.01$ $80.8\%$ $0.13\pm 0.02$ $73.1\%$
5.0.2 ZINC-250k
The experimental results of targeting logP and QED (alone and in conjunction) are presented in Table 2. Uniqueness and novelty scores are not reported as they approached 100% in almost all cases. Similarly to QM9, we observe that FreeGress outperforms DiGress on all tasks by a wide margin. For example, on the LogP task, FreeGress with $\rho=0.2,s=3$ achieves a 78% decrease in MAE with respect to the best DiGress variant ($\lambda=400$), while at the same time improving validity. On the QED task, FreeGress with $\rho=0.1,s=2$ achieves a 71% MAE decrease with respect to the best DiGress variant ($\lambda=600$), while keeping the same validity rates. When conditioning on LogP and QED in conjunction, FreeGress with $\rho=0.1,s=2$ decreases MAE by 65% in MAE compared to the best DiGress variant ($\lambda=800$) while slightly improving validity by 5%. Overall, the results empirically confirm that FreeGress performs better than DiGress regardless of the dataset where it is applied and the properties that is tasked to target.
Table 3: Results of applying the proposed node inference method $p_{\varepsilon}(n\,|\,\bm{y})$ to FreeGress while targeting Molecular Weight on the ZINC-250k dataset. The last column reports the improvement against a “no inference” FreeGress variant.
No inference $p_{\varepsilon}(n\,|\,\bm{y})$ Improvement
MAE $\downarrow$ Val. $\uparrow$ MAE $\downarrow$ Val. $\uparrow$ MAE $\downarrow$ Val. $\uparrow$
DiGress
[…]
$\lambda=400$ $64.75$ p m 6.05 79.3% $22.51$ p m 2.62 38.4% -65.2% -51.6%
$\lambda=600$ $64.67$ p m 6.17 80.2% $21.70$ p m 3.00 41.2% -66.4% -48.6%
$\lambda=800$ $62.96$ p m 6.21 77.8% $20.92$ p m 2.90 40.4% -66.8% -48.1%
FreeGress
$\rho=0$ $\bm{7.08\pm 1.19}$ 54.5% $7.25$ p m 1.81 75.7% +2.4% +38.9%
$\rho=0.1,s=1$ $14.93$ p m 2.27 60.9% $8.22$ p m 0.77 81.4% -44.9% +33.7%
$\rho=0.1,s=2$ $11.37$ p m 2.81 50.8% $8.96$ p m 1.93 79.7% -21.2% +56.9%
$\rho=0.1,s=3$ $9.33$ p m 1.87 46.5% $8.03$ p m 1.02 66.6% -13.9% +43.2%
$\rho=0.2,s=1$ $20.00$ p m 5.47 61.3% $13.22$ p m 1.51 81.3% -33.9% +32.6%
$\rho=0.2,s=2$ $13.69$ p m 3.19 53.9% $13.37$ p m 1.47 78.4% -2.3% +45.5%
$\rho=0.2,s=3$ $15.61$ p m 4.08 48.4% $14.28$ p m 2.12 75.0% -8.5% +54.9%
FreeGress+extra features
$\rho=0$ $12.66$ p m 4.99 60.1% $5.13$ p m 0.38 $\bm{86.1\%}$91.85% +43.25%
$\rho=0.1,s=1$ $43.27$ p m 9.35 82.1% $3.52$ p m 0.34 83.7% -91.87% +1.95%
$\rho=0.1,s=2$ $48.25$ p m 11.03 73.3% $\bm{3.34\pm 0.31}$ 84.7% -93.08% +15.55%
$\rho=0.1,s=3$ $47.67$ p m 13.26 63.2% $3.35$ p m 0.40 81.2% -92.97% +28.48%
5.0.3 Effect of node inference in ZINC-250k
Table 3 summarizes the results on applying node inference on the MW targeting task. The “no inference” scenario (first column) indicates that FreeGress obtains lower MAEs than DiGress, regardless of whether additional features are added or not. However, this happens at the expense of validity, which is lower for FreeGress than DiGress. Adding node inference (central column) allows FreeGress to generate valid molecules at the same rate than DiGress, but with better MAE. Moreover, the addition of node inference is beneficial to FreeGress itself, as quantified by the third column which measures row-wise improvement. Surprisingly, we observe that applying node inference to DiGress decreases MAE by approximately 71% across experiments, but at the same time completely disrupts chemical validity. In contrast, adding node inference to FreeGress causes a good decrease in MAE and a steep increase in validity (up to 57%). Overall, the best improvement is given by the combination of node inference and additional molecular features.
Input $\mu$: 0.0603
Est. $\mu$: 0.0463
Input $\mu$: 4.2338
Est. $\mu$: 4.1238
Input HOMO: -6.8083
Est. HOMO: -6.8604
Input HOMO: -7.4559
Est. HOMO: -7.4632
Input LogP: -0.9834
Est. LogP: -0.9246
Input QED: 0.5034
Est. QED: 0.5297
Input MW: 159.05
Est. MW: 155.03
Input MW: 401.13
Est. MW: 405.10
Figure 3: Curated molecules from the QM9 (top row) and ZINC-250k (bottom row). The input conditioning value and the one estimated (Est.) after generation are displayed below each molecule.
5.0.4 Curated samples.
In Figure 3 we show some curated samples from the targeting properties under study under study to add visual evidence that the model is operating correctly and it is able to target different properties with success. Additional random samples are shown in the Supplementary material.
@@ 6 Conclusions
We introduced FreeGress, a discrete DDPM for graphs with CF guidance, which allows to generate molecules complying with a pre-specified set of desired chemical properties. Through extensive experiments, we have shown that CF guidance allows to generate better (more tailored to the specification) molecules than CB guidance without sacrificing chemical validity. Additionally, we have implemented a form of learned node inference, and shown that inferring the number of nodes from the guide helps the generation whenever the molecular size is connected to the target property.
In future studies, we intend to tackle some of the current limitations of FreeGress to expand its applicability. A first research direction is to improve chemical validity designing the forward and reverse processes to use conditional transition matrices, thus constraining the intermediate noising and denoising steps to happen in valid chemical space. Another promising direction is to work with molecular fragments rather than atoms, as it would likely reduce the opportunities to generate an invalid intermediate structure. We intend to validate these findings also in the more general setting of arbitrary graph generation.
{credits}
6.0.1 Acknowledgements
Research partly funded by PNRR - M4C2 - Investimento 1.3, Partenariato Esteso PE00000013 - "FAIR - Future Artificial Intelligence Research" - Spoke 1 "Human-centered AI", funded by the European Commission under the NextGeneration EU programme.
[…]
The authors have no competing interests to declare that are relevant to the content of this article.